OpenStudy (anonymous):

Trigonometry with Unit Circles Question! Refer to the graph of y=sin(x) or y=cos(x) to find the exact values of x in the interval [0,4pi] that satisfy the equation. Sin(x) = 1/2 So I used Pi/6 [squarerootof3/2, 1/2] as my reference angle on the unit square. I went around the circle twice (up to 4pi) counter-clockwise (giving me Pi/6, 5pi/6, 13pi/6, and 17pi/6) and then twice clockwise (giving me -7pi/6, -11pi/6, -19pi/6, and -23pi/6). I entered these values in a comma separated form, but it is not accepted as the right answer. Can anyone see what I am missing?

OpenStudy (anonymous):

Well, since \(0 \le x \le 4\pi\) then the negative values you got are not apart of the solution. The answer should be only your positive answers (counter-clockwise), and as reference: http://www.wolframalpha.com/input/?i=Sin%28x%29+%3D+1%2F2%2C+0%3C%3Dx%3C%3D4pi

OpenStudy (jdoe0001):

hmmm I don't see anything wrong with your values but it may be likely that it's as pitamar is saying, that may be you're only expected to enter the positive angles